# Error-quantified Conformal Inference for Time Series

### v1

@@ 3.1 Error quantification via smoothing feedback
However, subgradients of quantile loss $\ell_{t}$ can only take two values $1-\alpha$ and $-\alpha$, regarding the feedback value of coverage ($s_{t}\leq q_{t}$) and miscoverage ($s_{t}>q_{t}$) respectively. As a consequence, the feedback keeps the same value no matter how severe the miscoverage is or how conservative the coverage is. In other words, discrete feedback value does not exploit the information of the error.
To address this issue, we consider smoothing the feedback when updating the thresholds. Let $f(x)\in[0,1]$ be a smooth approximation to the indicator $\mathds{1}(x>0)$, such as Sigmoid function, Gaussian error function, etc. Then we can approximate the quantile loss $\ell_{t}(q)$ via ~$\tilde{\ell}_{t}(q)=(f(s_{t}-q)-\alpha)(s_{t}-q)$. This smoothing technique was demonstrated in previous studies (Kaplan & Sun, 2017; Fernandes et al., 2021). Applying OGD on the smoothed quantile loss {~$\{\tilde{\ell}_{t}(q)\}_{t\geq 1}$, we have the following fully smoothed update rule
@@ 3.2 Extended versions
where $\{w_{i}\}_{1\leq i\leq t}\subseteq[0,1]\text{ is a sequence of increasing weights with }\sum_{i=1}^{t}w_{i}=1.$ Equation (7) evaluates the recent empirical miscoverage frequency and degree of miscovery when deciding whether or not to lower or raise $q_{t}$, making coverage more stable.
@@ 3.3 Distribution-free coverage guarantees
In this section, we outline the theoretical coverage guarantees of ECI. The detailed proofs are deferred in Appendix B. We first present two assumptions and briefly explain their feasibility.
@@
The following theorem provides a dynamic miscoverage bound when the learning rate is fixed. We prove it by showing that every miscoverage step will be followed by at least $N-1$ coverage step, where $N=\lfloor\alpha^{-1}\rfloor$, that is if $Y_{t}\notin\hat{C}_{t}$ happens, then $Y_{t+i}\in\hat{C}_{t+i}$ holds for $i=1,2,\cdots,N-1$.
@@ 4.1 Setup
General implement. We choose target coverage $1-\alpha=90\%$ and construct asymmetric prediction sets using two-side quantile scores under $\alpha/2$ respectively. For each baseline, we select the most appropriate range of learning rates $\eta$ for the respective datasets and present the best results in the tables. For sets, all baselines will output asymmetric sets $[\hat{Y}_{t}-q^{l}_{t},\hat{Y}_{t}+q^{u}_{t}]$ with upper score $q^{u}_{t}$ and lower score $q^{l}_{t}$ under half of the coverage level $\alpha/2$ respectively.
@@ 4.2 Results in finance domain
decay-OGD 90.1 0.531 0.521 90.1 0.122 0.099 90 0.100 0.059
PID 90.1 0.207 0.177 90 0.434 0.432 89.9 0.413 0.411
ECI 90 0.384 0.382 90 0.117 0.098 89.9 0.071 0.055
ECI-cutoff 90 0.405 0.396 90.2 0.118 0.096 90.1 0.072 0.055
ECI-integral 90.1 0.402 0.398 90 0.117 0.098 89.9 0.072 0.055
@@ 4.5 Results in synthetic data
ACI 89.9 $\infty$ 8.2 89.9 $\infty$ 8.2 89.9 $\infty$ 8.43
OGD 90 8.49 8.5 89.9 8.39 8.4 89.9 8.73 8.7
SF-OGD 90 12.48 11.56 90 12.58 11.69 89.9 12.7 11.88
decay-OGD 90 8.3 8.22 90 8.26 8.21 90 8.57 8.6
@@ A.2 Fully smoothed ways
Experimental results also demonstrate it. Since $f(x)$ does not approach 00 quickly in the early part of the period, and may even be larger than $\alpha$, the coverage rate can not effectively approach $1-\alpha$. It can be seen that smoothed method and fully smoothed method tends to have overly high coverage in the early stages in the Figures 5, 6 and 7. Consequently, it results in wider sets.
@@ B.2 Proof of Theorem 1
In conclusion, we have proved that for every miscoverage step $t$, i.e. $Y_{t}\notin\hat{C}_{t}$, the next $N-1$ steps of (5) will satisfy $Y_{t+i}\in\hat{C}_{t+i},i=1,2,\cdots N-1$. Therefore, for any $T$,
@@ C.2 OGD
$\alpha\in(0,1)$, base predictor $\hat{f}$, learning rate $\eta>0$, init. $q_{1}\in\mathbb{R}$
@@ C.3 SF-OGD
$\alpha\in(0,1)$, base predictor $\hat{f}$, learning rate $\eta>0$, init. $q_{1}\in\mathbb{R}$
@@
$\alpha\in(0,1)$, base predictor $\hat{f}$, learning rate $\eta>0$, init. $q_{1}\in\mathbb{R}$
@@ C.5 Conformal PID
$\alpha\in(0,1)$, base predictor $\hat{f}$, trained scorecastor $g^{\prime}$, $\eta>0$, window length $w$, init. $q_{1}\in\mathbb{R}$
@@ C.6 SPCI
5:     Use quantile regression to obtain $\hat{Q}_{t}\leftarrow\mathcal{Q}(\hat{\varepsilon})$
@@ Appendix E Experimental results in synthetic data
$\beta_{t}=\beta^{(0)}=(2,1,0,0)^{\top},\quad t=1,\ldots,500,$
$\beta_{t}=\beta^{(1)}=(0,-2,-1,0)^{\top},\ t=501,\ldots,1500,$
$\beta_{t}=\beta^{(2)}=(0,0,2,1)^{\top},\quad t=1501,\ldots,2000,$

### v2

@@ 3.1 Error quantification via smoothing feedback
However, subgradients of quantile loss $\ell_{t}$ can only take two values $1-\alpha$ and $\alpha$, regarding the feedback value of coverage ($s_{t}\leq q_{t}$) and miscoverage ($s_{t}>q_{t}$) respectively. As a consequence, the feedback keeps the same value no matter how severe the miscoverage is or how conservative the coverage is. In other words, discrete feedback value does not exploit the information of the error.
To address this issue, we consider smoothing the feedback when updating the thresholds. Let $f(x)\in[0,1]$ be a smooth approximation to the indicator $\mathds{1}(x>0)$, such as Sigmoid function, Gaussian error function, etc. Then we can approximate the quantile loss $\ell_{t}(q)$ via $\tilde{\ell}_{t}(q)=(f(s_{t}-q)-\alpha)(s_{t}-q)$. This smoothing technique was demonstrated in previous studies (Kaplan & Sun, 2017; Fernandes et al., 2021). Applying OGD on the smoothed quantile loss $\{\tilde{\ell}_{t}(q)\}_{t\geq 1}$, we have the following fully smoothed update rule
@@ 3.2 Extended versions
[…]
@@ 3.3 Distribution-free coverage guarantees
In this section, we outline the theoretical coverage guarantees of ECI. The detailed proofs are referred in Appendix B. We first present two assumptions and briefly explain their feasibility.
@@
The following theorem provides a dynamic miscoverage bound when the learning rate is fixed. We prove it by showing that every miscoverage step will be followed by at least N1N-1 coverage step, where $N=\lfloor\alpha^{-1}\rfloor$, that is if $Y_{t}\notin\hat{C}_{t}$ happens, then $Y_{t+i}\in\hat{C}_{t+i}$ holds for $i=1,2,\cdots,N-1$.
@@ 4.1 Setup
General implement. We choose target coverage $1-\alpha=90\%$ and construct asymmetric prediction sets using two-side quantile scores under $\alpha/2$ respectively. For each baseline, we select the most appropriate range of learning rates $\eta$ for the respective datasets and present the best results in the tables. For sets, all baselines will output asymmetric sets $\hat{Y}_{t}-q^{l}_{t},\hat{Y}_{t}+q^{u}_{t}]$ with upper score $q^{u}_{t}$ and lower score $q^{l}_{t}$ under half of the coverage level $\alpha/2$ respectively.
@@ 4.2 Results in finance domain
decay-OGD 90.1 0.531 0.521 90.1 0.122 0.099 90.0 0.100 0.059
PID 90.1 0.207 0.177 90.0 0.434 0.432 89.9 0.413 0.411
ECI 90.0 0.384 0.382 90.0 0.117 0.098 89.9 0.071 0.055
ECI-cutoff 90.0 0.405 0.396 90.2 0.118 0.096 90.1 0.072 0.055
ECI-integral 90.1 0.402 0.398 90.0 0.117 0.098 89.9 0.072 0.055
@@ 4.5 Results in synthetic data
ACI 89.9 $\infty$ 8.20 89.9 $\infty$ 8.20 89.9 $\infty$ 8.43
OGD 90.0 8.49 8.50 89.9 8.39 8.40 89.9 8.73 8.70
SF-OGD 90.0 12.48 11.56 90.0 12.58 11.69 89.9 12.70 11.88
decay-OGD 90.0 8.30 8.22 90.0 8.26 8.21 90.0 8.57 8.60
@@ A.2 Fully smoothed ways
Experimental results also demonstrate it. Since $f(x)$ does not approach 0 quickly in the early part of the period, and may even be larger than $\alpha$, the coverage rate can not effectively approach $1-\alpha$. It can be seen that smoothed method and fully smoothed method tends to have overly high coverage in the early stages in the Figures 5, 6 and 7. Consequently, it results in wider sets.
@@ B.2 Proof of Theorem 1
In conclusion, we have proved that for every miscoverage step $t$, i.e. $Y_{t}\notin\hat{C}_{t}$, the next N1N-1 steps of (5) will satisfy $Y_{t+i}\in\hat{C}_{t+i},i=1,2,\cdots N-1$. Therefore, for any $T$,
@@ C.2 OGD
[…]
@@ C.3 SF-OGD
[…]
@@
[…]
@@ C.5 Conformal PID
[…]
@@ C.6 SPCI
[…]
@@ Appendix E Experimental results in synthetic data
[…]